OpenStudy (anonymous):

Find the equation of the line parallel to y = -4x - 3 and passing through the point (-2, 7)

OpenStudy (anonymous):

With parallel lines, the slope is the exact same which means for our new equation our slope will be -4. Now all we have to do is put this into point-slope. y-y1=m(x-x1) [Now we will pluck the coords into the equation] y-7=-4(x-(-2)) [The --2 becomes +2 and we distribute] y-7=-4(x+2) y-7=-4x-8 [Now we add 7 to both sides and get] y=-4x-1